### v4

# Shamier Settle, Matthew Shapiro, Emerita Torres, and Anand Vidyashankar for discussion and feedback on the work described

Abstract
We estimate the number of street vendors in New York City. First, we summarize the process by which vendors receive licenses and permits to operate legally in New York City. Second, we describe a survey that was administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit. Third, we review ratio estimation and provide a theoretical justification based on the theory of point processes. Fourth, we use ratio estimation to calculate the total number of vendors, finding approximately 23,000 street vendors operate in New York City—20,500 mobile food vendors and 2,400 general merchandise vendors—with one third located in just six ZIP Codes—11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. Finally, we evaluate the accuracy of the ratio estimator when the distribution of vendors is explained by a Poisson or Yule process, and we discuss several policy implications. In particular, our estimates suggest the American Community Survey misses the majority of New York City street vendors.
1. Introduction
Street vendors are New York City’s smallest businesses, selling food and merchandise from carts, stalls, and trucks throughout the five boroughs. They are an iconic part of the urban landscape and a thriving sector of the local economy, contributing millions of dollars in government revenue through taxes, fines, and fees. [^1] Fees from mobile food vending licenses and permits provide nearly a million dollars in revenue to New York City each year (Mosher and Turnquist 2024). Fines paid by vendors provide approximately \$200,000 each year according to data from the New York City Office of Administrative Trials and Hearings. Perhaps most importantly, street vending historically benefits underserved communities, both because vendors operate in neighborhoods with limited access to traditional stores and because vending is one of a handful of occupations in which New Yorkers of all backgrounds, immigrants in particular, can achieve economic mobility and a chance at the American dream (Burrows and Wallace 1998, chap. 42).
Yet despite their social and economic importance, little is known about the size and location of New York City’s street vending population. This is because while local law requires street vendors to obtain licenses and permits to operate legally, the number of licenses and permits are limited, resulting in a largely unknown population of vendors that operate without a license or permit. These vendors are not easily identified from administrative datasets, such as tax records or fines, and they can be difficult to locate for government surveys, such as the American Community Survey. Nevertheless, understanding the size and location of New York’s street vending population is crucial for informing policy and advocacy.
In this paper, we propose an estimate for the number of street vendors in New York City, including those that operate without a license or permit. We present our work in four sections. In Section 2, we review the process by which vendors receive licenses and permits, and we describe a survey administered by the Street Vendor Project at the Urban Justice Center while distributing Coronavirus relief aid to vendors operating both with and without a license or permit.
In Section 3, we use the fact that the number of licenses and permits are limited by law to construct a ratio estimator for the number of street vendors. We then provide a theoretical justification of our estimates, assuming the spatial distribution of survey respondents is representative and well-approximated by a set of inhomogeneous Poisson processes. We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors. One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. We also derive standard errors.
In Section 4, we examine the validity of our assumptions. We evaluate whether the spatial distribution of respondents is representative by comparing it to the American Community Survey and administrative data, and we evaluate the Poisson process assumption by considering a more general model in which the arrival of street vendors follows a pure birth process, of which the Poisson and Yule processes are special cases. The Yule process may be appropriate if street vendors cluster within markets, in which case the data might have extra-Poisson variation. We rederive the standard errors from Section 3 to account for this variation.
In Section 5, we discuss some implications of our estimates. In particular, we find that though the spatial distribution of respondents matches the results of the American Community Survey, the two disagree on the size of New York City’s street vending population. Our estimates suggest the American Community Survey misses the majority of New York City street vendors.
2. Background and Survey Design
A street vendor is any individual who sells goods from a mobile vending unit instead of a store. We distinguish between the mobile vending unit from which the goods are sold (i.e., the vending establishment) and the individuals who own and/or operate that unit (i.e., the vendors). Note that an establishment refers to a single vending unit. There may be multiple vendors associated with any establishment, and multiple establishments may be owned by a single individual or firm. [^2] For example, suppose Sally owns the business Sally’s Salads, which consists of two food trucks, each staffed by three employees. Then Sally’s Salads counts as one business, two establishments, and seven vendors.
We limit our analysis to two types of street vendors: mobile food vendors—vendors who sell food items such as sandwiches, drinks, and cut fruit—and general merchandise vendors—vendors who sell merchandise items such as electronics, clothing, and accessories. We refer to these vendors as food vendors and merchandise vendors, respectively. In Section 2.1, we review the process by which vendors receive licenses and permits as required by law. In Section 2.2, we describe a survey administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit.
2.1 Background
A long list of rules determines the individuals, locations, and time periods during which vendors can legally sell their goods in New York City. Most relevant is the fact that a merchandise vendor requires a license to operate legally, and a food vendor requires both a license and a permit. The number of food permits and merchandise licenses are limited by law: 5,100 permits and 853 general merchandise licenses are available. Mobile food vending licenses are not limited. Note that a mobile food vending permit is issued to an individual or business to allow for the sale of food from a specific mobile food vending unit, such as a cart or truck. A mobile food vending license authorizes an individual to prepare or serve food from a permitted mobile food vending unit. [^3] Note that our estimate reflects the number of street vendors as of 2021, prior to the implementation of Local Law 18 of 2021, which was enacted concurrently. Local Law 18 made several changes to the license and permit process. These changes do not impact our estimates and therefore are not discussed in this paper.
There are several types of mobile food vendor permits with varying restrictions. Of the 5,100 permits available, 200 are borough permits that limit vendors to one of the five boroughs; 100 are reserved for veterans or vendors with a disability; 1,000 are seasonal and valid only from April to October; 1,000 are green cart permits that limit vendors to selling fruit, vegetables, plain nuts, and water; and 2,800 are unrestricted. Multiple individuals with food vendor licenses can legally operate from one permitted food vending unit. Merchandise licenses are renewed annually, while food licenses and permits are renewed biennially.
There are three relevant exceptions to these rules. The first exception is that merchandise vendors who are veterans are not subject to the limit of 853 licenses. According to data obtained by Mosher and Turnquist (2024), there are approximately 1,000 licensed merchandise vendors who are veterans in New York City. We assume the number of veterans selling merchandise without a license is negligible.
The second exception is First Amendment vendors. A First Amendment vendor sells expressive merchandise such as newspapers, books, and art. Expressive merchandise is considered free speech and protected under the First Amendment of the U.S. Constitution so that the number of First Amendment vendors cannot be restricted by law.
The final exception are concessionaires that operate on New York City parkland through an NYC Parks license or permit. According to the New York City Department of Parks and Recreation, there are approximately 400 concessions within New York City parks. Many offer food services, ranging from food carts to restaurants. We do not consider First Amendment vendors or park concessionaires in this paper. Our estimates are limited to food and merchandise vendors operating outside of parks.
2.2 Survey Design
The Street Vendor Project (SVP), part of the Urban Justice Center, is a non-profit organization that advocates on behalf of New York City street vendors. SVP administered a survey to approximately 2,000 street vendors while distributing Coronavirus relief aid in 2021. The aid was a one-time payment of \$1,000, available to any individual that owned, operated, or was otherwise employed by a street vending business in New York City between 2020 and 2021. First Amendment vendors were also eligible. There were no limits based on the size of the business, the number of sales, or whether licensed and/or permitted. All individuals eligible for aid were invited to complete a survey.
The population of street vendors estimated in this paper is the population eligible for aid (as determined by SVP) that self-identifies as either a food or merchandise vendor. SVP found aid-eligible individuals through membership lists, referrals, and canvassing operations in which SVP affiliates visited neighborhoods. Survey operations continued until \$2,415,000 in funds were distributed, yielding 2,060 responses.
The survey included 100 questions and was conducted in nine different languages: Arabic, Bangla, Cantonese, English, French, Mandarin, Spanish, Tibetan, and Wolof. [^4] Vendors who spoke other languages, such as Dari, Farsi, Russian, and Turkish, typically spoke English with a high level of fluency, and those surveys were conducted in English. Similarly, vendors who spoke indigenous languages from Latin America, such as Nahuatl and Quechua, typically spoke Spanish with a high level of fluency. The survey items solicited a variety of information from vendors, such as logistical information (e.g., vending location, residential location, and frequency of operation), economic information (e.g., items sold, income, and expenses), and demographic information (e.g., age, race, ethnicity, and immigration status). Most relevant is the fact that respondents classified themselves by the goods they sold (e.g., food vendors, merchandise vendors, First Amendment vendors, etc.), and respondents indicated whether they had the relevant licenses and permits to vend. Of the 2,060 responses, 1,400 identified as food vendors and 559 as merchandise vendors. The remaining 101 respondents were predominantly First Amendment Vendors, which we exclude from our analysis.
Of the 1,400 food vendors, 349 (25%) indicated they had a permit to vend. Of the 559 merchandise vendors, 505 were not veterans, of which 308 (61%) indicated they had a license to vend. The number of respondents is listed by neighborhood in Section 7.1, Table 1. (Neighborhoods with few respondents are grouped together.) A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Section 7.2, Figure 1. Approximately five percent of vendors did not identify their vending location. These vendors are included in the New York City total, but they are excluded from the neighborhood estimates in Table 1 and Figure 1. (For this reason, the Respondent and Population columns do not sum exactly to the New York City total.) Veteran merchandise vendors are excluded entirely from Table 1 and Figure 1, although these vendors are reflected in our overall estimate of 23,000 vendors.
3. Methodology and Data Analysis
We estimate approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors. We arrive at the estimate by using ratio estimation, which leverages the fact that the number of licenses and permits are limited by local law.
In Section 3.1, we provide a simple explanation of ratio estimation, which is intended to be accessible to a general readership. In Section 3.2, we provide a theoretical justification of our estimates based on the theory of point processes, which, while more technical, is intended to provide a somewhat general argument for ratio estimation with spatial data when a design-based approach cannot be justified. Inference is conducted using increasing domain asymptotics, and the details are outlined in Section 7.3. In Section 3.3, we apply the methodology to the survey data discussed in Section 2.2.
3.1 Ratio Estimation
Ratio estimation is a common approach for estimating the size of a population. See Cochran (1978) and Hald (1998, chap. 16) for a historical discussion and Lohr (2021, chap. 4) for an introduction. We provide a simple explanation based on cross-multiplication, also called the rule of three. The purpose is to highlight the main assumption. A more rigorous derivation is provided in Section 3.2.
Consider a fixed region $A$, and let $N_{i}(A)$ denote the number of individuals in region $A$ of type $i$. To fix ideas, let $A$ be New York City, let $i=0$ indicate mobile food vendors with a permit, and let $i=1$ indicate mobile food vendors without a permit. Also, let $n_{i}(A)$ denote the number of respondents of type $i$.
We observe $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$ from which we estimate $\theta(A)=\mathbb{E}\left[N_{0}(A)\right]$, the expected number of vendors without a permit. The expected number of mobile food vendors with or without a permit is then $\tau(A)=\theta(A)+N_{1}(A)$.
The ratio estimator of $\theta(A)$ is
$\hat{\theta}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
It is called the ratio estimator because it depends on the ratio of random variables $n_{0}(A)\,/\,n_{1}(A)$. (The total number of permitted vendors, $N_{1}(A)$, is considered fixed.) The estimator is accurate when the response rate, $p$, is approximately the same for each type, i.e.,
$p\approx\frac{n_{0}(A)}{N_{0}(A)}\approx\frac{n_{1}(A)}{N_{1}(A)}$
in which case $\hat{\theta}(A)$ is the solution for $N_{0}(A)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$, obtained by cross-multiplication.
The main assumption of ratio estimation is that the survey data are representative in the sense that the response rate does not depend on whether a mobile food vendor has a permit. If we further assume that the response rate does not vary by subregion, we can also estimate the total number of mobile food vendors by subregion. For example, suppose the response rate in subregion $B\subset A$ is also approximately equal to $p$, e.g.,
$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{0}(B)}{N_{0}(B)}$
Then we can estimate $\theta(B)$ by the subregion estimator
$\tilde{\theta}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
where $\tilde{\theta}(B)$ is the solution for $N_{0}(B)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(B)$.
The subregion estimator is necessary when $N_{1}(B)$ is not observed and $\hat{\theta}(B)$ cannot be calculated directly. This is the case with the survey data described in Section 2.2. We observe $N_{1}(A)$, the number of permits city-wide (i.e., within region $A$), but not $N_{1}(B)$, the number of permits within subregion $B$. (We also observe $n_{0}(B)$ and $n_{1}(B)$.)
The subregion estimator $\tilde{\theta}(B)$ works by replacing the term $N_{1}(B)$ in the ratio estimator $\hat{\theta}(B)$ with the estimate
$\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,n_{1}(B)}{n_{1}(A)}$
The estimate $\tilde{N}_{1}(B)$ is also used to estimate the total number of mobile food vendors expected in subregion $B$. That is, we estimate $\tau(B)=\theta(B)+N_{1}(B)$ with
$\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}$
3.2 Model
We provide a point process justification of the ratio estimation procedure described in Section 3.1. The purpose is to more closely examine the underlying assumptions and derive the standard errors of the estimates.
Let $\Pi_{i}$ denote an inhomogeneous spatial Poisson process referencing the location $x\in A\subset\mathbb{R}^{2}$ of each street vendor with status $i$ at the time the survey was conducted. As in Section 3.1, we fix ideas by letting $i=0$ indicate mobile food vendors with a permit and $i=1$ indicate mobile food vendors without a permit.
The Poisson process assumption may be justified by the law of rare events. New York City can be partitioned into a large number of theoretically vendable locations. Whether a vendor is located within a partition has a vanishingly small probability, such that the number of vendors in any area is well approximated by a Poisson distribution. This approximation is accurate even if these probabilities are weakly dependent, see Freedman (1974). Other justifications are possible. For example, see Section 4.2 for a justification of the Poisson process based on the theory of birth processes.
[…]
In Section 3.1, we described the main assumption of ratio estimation: that the status and spatial distribution of the respondents is representative. We now make this statement precise. We assume that $p_{i}(x)$ can be decomposed into a constant that does not depend on $i$ or $x$ plus a spatially varying error term, $p_{i}(x)=p\,+\,\epsilon_{i}(x)$, and the error term is orthogonal to the corresponding mean measure. i.e.,
[…]
It follows that
[…]
so that by conditioning on $N_{1}(B)$, we arrive at the following probability model for the number of respondents with and without permits
$n_{0}(B)\sim$ $\ \text{Poisson}\big{(}p\,\theta(B)\big{)}$
$n_{1}(B)\ |\ N_{1}(B)\sim$ $\ \text{Binomial}\big{(}p,\,N_{1}(B)\big{)}$
The maximum likelihood estimates for $p$ and $\theta(A)$ are
$\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}\qquad\text{and}\qquad\hat{\theta}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
Using increasing domain asymptotics, the estimator is asymptotically normal with mean $\theta(A)$ and standard error
$\text{SE}[\hat{\theta}(A)]=\theta(A)\sqrt{\dfrac{1}{p\,\theta(A)}+\dfrac{1-p}{p\,N_{1}(A)}}$
Substituting $\hat{p}$ and $\hat{\theta}(A)$ for $p$ and $\theta(A)$ yields the following plug-in estimator for the standard error,
$\hat{\text{SE}}[\hat{\theta}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
See Section 7.3.2 for details.
The maximum likelihood estimate for $\theta(B)$ is
$\tilde{\theta}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
which is asymptotically normal with mean $\theta(B)$ and standard error
$\text{SE}[\tilde{\theta}(B)]=\theta(B)\sqrt{\frac{1}{p\,\theta(B)}+\frac{1-p}{p\,N_{1}(A)}}$
The plug-in estimator for the standard error is
$\tilde{\text{SE}}[\tilde{\theta}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
To determine the subregion total $\tau(B)=\theta(B)+N_{1}(B)$, we estimate $\theta(B)$ and $N_{1}(B)$ with $\tilde{\theta}(B)$ and $\tilde{N}_{1}(B)=N_{1}(A)\,n_{1}(B)\,/\,n_{1}(A)$, yielding
$\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(B)+n_{1}(A\backslash B)}$
where $A\backslash B$ denotes the compliment of subregion $B$.
The asymptotic standard error is
$\text{SE}[\tilde{\tau}(B)]=\theta(B)\sqrt{\frac{1}{p\,\theta(B)}+\frac{1-p}{p\,N_{1}(A)}\left(1+\frac{N_{1}(B)\,N_{1}(A\backslash B)}{\theta(B)^{2}}\right)}$
and the plug-in estimator is
$\tilde{\text{SE}}[\tilde{\tau}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\left(\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}\right)\left(1+\dfrac{n_{1}(B)\,n_{1}(A\backslash B)}{n_{0}(B)^{2}}\right)}$
See Section 7.3.3 for details.
3.3 Data Analysis
We use the methodology described in Section 3.1 and 3.2 to estimate the number of street vendors in New York City. According to the survey data described in Section 2.2, $n_{1}(A)=349$ mobile food vendor respondents indicated they had a permit, and $n_{0}(A)=1,051$ indicated they did not. Since there are $N_{1}(A)=5,100$ permits available, it follows that the number of vendors without permits is 15,3505,100×1,051/ 349formulae-sequence153505100105134915,350$\approx$ 5,100$\times$ 1,051\,/\,34915 , 350 ≈ 5 , 100 × 1 , 051 / 349. Combined with the $5,100$ permitted vendors, the total number of mobile food vendors is approximately $20,500$.
By similar reasoning—i.e., exploiting the fact that the number of general merchandise licenses is limited for non-veterans—the number of non-veteran merchandise vendors is estimated to be approximately $1,400$. When combined with the estimated $1,000$ veteran merchandise vendors (see Section 2.1) and $20,500$ food vendors, the total number of street vendors is approximately $23,000$.
Estimates for select subregions are listed in Section 7.1, Table 1. We find that approximately one quarter of street vendors are located in the Manhattan neighborhoods of Chelsea, Clinton, and Lower Manhattan. Another quarter is located in West and North Queens. We also report the margin of error (i.e., two standard errors, half of the width of a 95% confidence interval).
We check these estimates by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of street vendors was documented through on-the-ground observations. The independent assessment identified 188 total street vendors near Fordham Road (Fordham Road BID and Street Vendor Project 2024). According to the survey data, 17 respondents indicated that they operate near Fordham Road, yielding an estimate of 152 street vendors total with a standard error of 54. We conclude this estimate is consistent with the 188 street vendors counted independently by Fordham Road BID and Street Vendor Project (2024).
4. Model Validation
We make two assumptions in Section 3. The main assumption is that the status and spatial distribution of the respondents is representative. This assumption is necessary to ensure the ratio estimator is accurate in large samples (i.e., consistent). A secondary assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous spatial Poisson processes. This assumption is necessary to derive the standard errors.
In this section, we examine each assumption. Since neither can be evaluated from the survey data alone, we consider additional information. In Section 4.1, we compare the spatial distribution of the respondents to data from the New York City Office of Administrative Trials and Hearings and the American Community Survey. In Section 4.2, we model the arrival of street vendors as a pure birth process, and we derive a more general formula for the standard errors in which the amount of extra-Poisson variation can be determined by the number of vendors that cluster within markets. [^5] The results of this section can also be derived from a spatial Negative Binomial Lévy process in a manner that mirrors Section 3.2. The advantage of the birth process approach taken in Section 4.2 is that it provides a plausible explanation for why the Poisson assumption made in Section 3.2 may not hold in practice.
4.1 Representativeness
The main assumption in Section 3 is that the status and spatial distribution of the respondents is representative (i.e., the response probability is spatially uncorrelated with the mean measure). If this assumption holds, then the ratio estimator is consistent for a wider class of point processes in which Campbell’s Theorem applies, not just the Poisson process. See Daley and Vere-Jones (2007) for a general statement of Campbell’s Theorem.
If this assumption is violated, then $\hat{\theta}(A)$ may be inconsistent. For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(A)\,/\,n_{1}(A)$ would not be representative, and $\hat{\theta}(A)$ would likely underestimate $\theta(A)$ even in large samples.
We think the representativeness assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid. We argue that trust in SVP and its mission, along with the aid, removed many barriers that typically prevent survey operations from reaching hard-to-reach populations. Instead, nonresponse reflects chance variation in the personal circumstances of a vendor that are largely unrelated to location.
Nevertheless, it is possible that canvassing operations systematically missed locations, and these missed locations may be revealed by examining other data sources. We thus compare the spatial distribution of vending locations reported by respondents to the spatial distribution of vendors who allegedly violated street vending laws during the year 2021. We consider the location of an individual’s first violation in 2021, reported in administrative data from the NYC Office of Administrative Trials and Hearings (OATH) (accessed 2024-04-03). We also compare the residencies reported by the respondents to the residencies of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers (Standard Occupational Classification 41-9091) estimated in the 2017-2021 American Community Survey (ACS) Public Use Micro Sample from the U.S. Census Bureau (accessed 2022-12-22).
These comparisons are visualized in Section 7.2, Figure 1. We find the spatial distribution of vending locations largely agree (top panels) as do the spatial distribution of the residencies (bottom panels), although there are several small discrepancies. For example, the OATH data suggest a larger percentage of vendors work in South Brooklyn (Coney Island and Brighton Beach) than captured by the Street Vendor Project survey, and the ACS data suggest a larger percentage of street vendors live in South Queens (Rockaway). However, both the OATH and ACS data suggest these locations reflect only a small percentage of New York City vendors, and thus any resulting undercount is likely to have little impact on our overall estimate.
We also note that even if this discrepancy suggests a location was missed by the Street Vendor Project, our main assumption is not necessarily violated. That is because the undercount in some areas may be compensated by an overcount in others (e.g., the Upper East Side of Manhattan or East Brooklyn) such that our overall estimate is still accurate.
4.2 Overdispersion
The second assumption in Section 3.2 is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes. We consider this assumption secondary because the consistency of $\hat{\theta}(A)$ holds under a wide class of point processes described in Section 4.1. That said, the standard errors derived in Section 3.2 do not necessarily hold when the Poisson assumption is violated. [^6] The ratio estimator may no longer be asymptotically efficient as well. In Section 7.3, we show that under the Poisson process assumption, the ratio estimator maximizes the likelihood, which implies it is asymptotically efficient. However, the ratio estimator may not maximize the likelihood when the Poisson process assumption is violated.
We think the Poisson assumption is reasonable because a wide class of data generating processes are well-approximated by the Poisson process under the law of rare events as discussed in Section 3.2. Nevertheless, to evaluate how the standard errors might change when this assumption is violated, we consider a general model that describes how street vendors arrived at their present locations. Specifically, we consider two families of pure birth processes: one with a constant arrival rate (or “birth rate”) in which the Poisson assumption holds and one with a linear arrival rate in which the Poisson assumption does not hold.
[…]
Now suppose instead that larger markets grow faster than smaller markets, an empirical phenomenon known as preferential attachment. Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate, or Yule process (Yule 1925). That is, let $N^{t}_{i}(B)$ denote the number of vendors in subregion $B$ at time $t$ so that the time between vendor arrivals at each market follows an exponential distribution with rate $\lambda_{i}\,N^{t}_{i}(B)$. If we assume further that each market started with one vendor at $t=0$, then the number of vendors at the time the survey was conducted ($t=1$), $N_{i}(B)$, would follow a negative binomial distribution with mean $\mu_{i}(B)=\exp(\lambda_{i})\,M(B)$ and variance $\mu_{i}(B)\bigr{(}\mu_{i}(B)-M(B)\bigr{)}\,/\,M(B)$. i.e.,
$N_{i}(B)\sim\ \text{Negative Binomial}\left(\mu_{i}(B),\ \mu_{i}(B)\,\frac{\mu_{i}(B)-M(B)}{M(B)}\right)$
where we parameterize the negative binomial distribution by its mean and variance. In this case, $N_{i}(B)$ exhibits extra-Poisson variation, and the results of Section 3.2 do not hold. See Ross (2014, chap. 6) for details. [^7] Note that we define the negative binomial and negative hypergeometric distributions as the total number of trials required to get a predetermined number of successes. The negative binomial assumes sampling with replacement, while the negative hypergeometric distribution assumes sampling without replacement.
To derive the standard errors, suppose a portion of the markets, $p$, were selected for the survey so that
$n_{i}(B)\sim\ \text{Negative Binomial}\left(p\,\mu_{i}(B),\ p\,\mu_{i}(B)\,\frac{\mu_{i}(B)-M(B)}{M(B)}\right)$
The conditional distribution of $n_{1}(B)$ given $N_{1}(B)$ is then
$n_{1}(B)\ |\ N_{1}(B)\sim\ \text{Negative Hypergeometric}\bigl{(}N_{1}(B)-1,\ M(B)-1,\ p\,M(B)\bigl{)}$
with conditional mean and variance
$\mathbb{E}\left[n_{1}(B)\ |\ N_{1}(B)\right]=p\,N_{1}(B)\qquad\text{and}\qquad\mathbb{V}\bigr{(}n_{1}(B)\ |\ N_{1}(B)\bigr{)}=w_{1}\,p\,(1-p)\,N_{1}(B)$
so that under increasing domain asymptotics, the ratio estimator
$\hat{\theta}(A)=\frac{N_{1}(A)\ n_{0}(A)}{n_{1}(A)}$
is asymptotically normal with mean $\theta(A)=\mu_{0}(A)$ and standard error
$\text{SE}[\hat{\theta}(A)]=\theta(A)\sqrt{w_{0}\left(\frac{1}{p\,\theta(A)}\right)+w_{1}\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
where
$w_{0}=\dfrac{\theta(A)-M(A)}{M(A)}\qquad\text{and}\qquad w_{1}=\dfrac{N_{1}(A)-M(A)}{M(A)+1}$
See Section 7.3.4 for details. Note that the weights $w_{0}$ and $w_{1}$ depend on $A$, although this dependence is not reflected in the notation.
The standard error derived in this section for $\hat{\theta}(A)$ is identical to the standard error for $\hat{\theta}(A)$ from Section 3.2 except for the weights $w_{0}$ and $w_{1}$. These weights are the result of the extra-Poisson variation induced by preferential attachment—that is, the linear “birth” rate in which the arrival of new vendors is proportional to the number of vendors in the market.
The weights $w_{0}$ and $w_{1}$ correspond to the relative number of vendors without permits and with permits per market respectively. These weights cannot be estimated from the survey data alone. However, we believe that $\theta(A)$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$. It follows that $w_{0}$ and $w_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most 3 times larger if the Poisson assumption is violated and the Yule process holds. We suspect this bound is conservative, however, and the typical market contains only one or two vendors.
The same behavior is exhibited by the subregion estimator of $\theta(B)$,
$\tilde{\theta}(B)=\dfrac{N_{1}(A)\ n_{0}(B)}{n_{1}(A)}$
which is asymptotically normal with mean $\theta(B)=\mu_{0}(B)$ and standard error
$\text{SE}[\tilde{\theta}(B)]=\theta(B)\sqrt{v_{0}\left(\frac{1}{p\,\theta(B)}\right)+w_{1}\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
where
$v_{0}=\dfrac{\theta(B)-M(B)}{M(B)}$
The estimator for the subregion total, $\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)$, is also asymptotically normal with mean $\tau(B)$ and standard error
$\text{SE}[\tilde{\tau}(B)]=\theta(B)\sqrt{v_{0}\left(\frac{1}{p\,\theta(B)}\right)+\left(\frac{v_{1}\,N_{1}(B)\left(N_{1}(A\backslash B)-\theta(B)\right)^{2}+v_{2}\,N_{1}(A\backslash B)\left(N_{1}(B)+\theta(B)\right)^{2}}{N_{1}(A)\,\theta(B)^{2}}\right)\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
where
$v_{1}=\dfrac{N_{1}(B)-M(B)}{M(B)+1}\qquad\text{and}\qquad v_{2}=\dfrac{N_{1}(A\backslash B)-M(A\backslash B)}{M(A\backslash B)+1}$
See Section 7.3.4 for details.
5. Discussion
We conclude by discussing the implications of our estimates. In particular, we assess the completeness of the New York City Office of Administrative Trials and Hearings (OATH) and American Community Survey (ACS) data. Our estimates suggest that while both data sets are spatially representative, they miss the majority of street vendors. We provide several explanations for this discrepancy.
We estimate 23,000 vendors work in New York City. In comparison, the 2017-2021 ACS estimates 4,634 individuals work the occupation of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in New York City. This suggests that, according to the ACS, there are at most 4,634 street vendors in New York City, and thus our finding indicates that the ACS estimate misses 80% of street vendors or more.
A fairer comparison might account for the fact that the standard error of our estimate is approximately 1,000, and the standard error of the ACS estimate is approximately 488, which we obtained from the replicate weights. Using the lower limit of a 95% confidence interval for our estimate (21,000) and the upper 95% confidence limit for the ACS estimate (5,610), our findings suggest the ACS misses roughly three quarters of street vendors or more.
Note the importance of the assumptions stated in Section 3.2 since other derivations of ratio estimation may imply smaller or larger standard errors, from which we might conclude the ACS misses a smaller or larger portion of the street vending population. For example, if the standard errors from Section 4.2 hold, then the confidence intervals would overlap if there were more than 300 vendors in the average market. We believe this is highly unlikely, suggesting the two estimates are inconsistent even under overdispersion.
Additional evidence comes from the OATH data, which indicate approximately 2,500 (unique) vendors are summoned to court each year for violating vending laws. If the ACS estimate holds, then more than half of vendors are summoned to court each year. If our estimate holds, then approximately one in ten vendors are summoned to court each year. We believe the latter is more realistic.
There may be several reasons why our estimates are significantly larger than the ACS. The fact that the Street Vendor Project distributed relief aid may have encouraged respondents who are unlikely to respond to the ACS. Moreover, the Street Vendor Project, through membership lists, referrals, and canvassing may have covered individuals missed by the Census Bureau’s sampling frame.
Another explanation is that the Street Vendor Project survey and the ACS may be measuring different populations or concepts. For example, the ACS estimate reflects the average number of vendors between 2017 and 2021, while our estimate reflects 2020 and 2021. For another example, the Census Bureau determines the occupation of an ACS respondent by autocoding responses to write-in questions. It is possible that the Street Vendor Project’s definition of a vendor is more inclusive. Street vending often provides supplemental income, and thus the discrepancy may reflect the broader challenge of studying the gig economy. In this case, either estimate may be relevant, depending on the intended use case.