### v5

# Shamier Settle, Matthew Shapiro, Emerita Torres, and Anand Vidyashankar for discussion and feedback on the work described

Abstract
We estimate the number of street vendors in New York City. We first summarize the process by which vendors receive licenses and permits to legally operate in New York City. We then describe a survey that was administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit. Finally, we calculate the total number of vendors using ratio estimation. We find that approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,300 general merchandise vendors. One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. We also provide a theoretical justification of our estimates based on the theory of point processes and a discussion of their accuracy and implications. In particular, our estimates suggest the American Community Survey fails to cover the majority of New York City street vendors.
1. Introduction
Street vendors are New York City’s smallest businesses, selling food and merchandise from carts, stalls, and trucks throughout the five boroughs. They are an iconic part of the urban landscape and a thriving sector of the local economy, contributing millions of dollars in government revenue through taxes, fines, and fees. Perhaps most importantly, street vending historically benefits underserved communities, both because vendors operate in neighborhoods with limited access to traditional stores and because vending is one of a handful of occupations in which New Yorkers of all backgrounds, immigrants in particular, can achieve economic mobility and a chance at the American dream (Burrows and Wallace 1998, chap. 42).
Yet despite their social and economic importance, little is known about the size and location of New York City’s street vending population. This is because while local law requires street vendors to obtain licenses and permits to operate legally, the number of licenses and permits are limited, resulting in a largely unknown population of vendors that operate without a license or permit. These vendors are not easily identified from administrative datasets, such as tax records or fines, and they can be difficult to locate for government surveys, such as the American Community Survey. Nevertheless, understanding the size and location of New York’s street vending population is crucial for informing policy and advocacy.
In this paper, we propose and justify an estimator for the number of street vendors in New York City, including those that operate without a license or permit. We present our work in six sections:
In Section 2, we review the process by which vendors receive licenses and permits, and we describe a survey administered by the Street Vendor Project at the Urban Justice Center while distributing Coronavirus relief aid to vendors operating both with and without a license or permit.
In Section 3, we use the fact that the number of licenses and permits are limited by law to construct a ratio estimator for the number of street vendors. We then develop a theoretical justification based on the theory of point processes, which assumes the spatial distribution of survey respondents is representative and well-approximated by a set of inhomogeneous Poisson processes. Finally, we derive the standard errors.
In Section 4, we examine the sensitivity of the ratio estimator to deviations from the underlying assumptions. We first consider the case in which the spatial distribution of the survey respondents is not representative, and we examine a more general estimator in which sampling weights adjust for differential nonresponse. We then consider the case in which the Poisson process assumption is violated, and we examine a more general model in which the arrival of street vendors follows a pure birth process, of which the Poisson and Yule processes are special cases. The Yule process may be more appropriate if street vendors cluster within markets. In both cases, we rederive the standard errors.
In Section 5, we estimate the number of street vendors in New York City. We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors. We also find that this estimate is somewhat insensitive to deviations from our underlying assumptions. For example, we find that if the spatial distribution of survey respondents is not representative, the estimated number of street vendors in New York City could be as low as 21,000 and as high as 32,000.
In Section 6, we discuss some implications of our estimates. In particular, we find that although the spatial distribution of respondents matches the results of the American Community Survey, the two disagree on the size of New York City’s street vending population. Our estimates suggest the American Community Survey misses the majority of New York City street vendors. This finding is important because the American Community Survey is commonly used for policy and advocacy, and our estimates shed light on the broader challenges of measuring the “gig economy.”
2. Background and Survey Design
A street vendor is any individual who sells goods from a mobile vending unit instead of a store. We distinguish between the mobile vending unit from which the goods are sold (i.e., the vending establishment) and the individuals who own and/or operate that unit (i.e., the vendors). Note that an establishment refers to a single vending unit. There may be multiple vendors associated with any establishment, and multiple establishments may be owned by a single individual or firm. For example, suppose Sally owns the business Sally’s Salads, which consists of two food trucks, each staffed by three employees. Then Sally’s Salads counts as one business, two establishments, and seven vendors.
We limit our analysis to two types of street vendors: mobile food vendors—vendors who sell food items such as sandwiches, drinks, and cut fruit—and general merchandise vendors—vendors who sell merchandise items such as electronics, clothing, and accessories. We refer to these vendors as food vendors and merchandise vendors, respectively. In Section 2.1, we review the process by which vendors receive licenses and permits as required by law. In Section 2.2, we describe a survey administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit.
2.1 Background
A long list of rules determines the individuals, locations, and time periods during which vendors can legally sell their goods in New York City. Most relevant is the fact that a merchandise vendor requires a license to operate legally, and a food vendor requires both a license and a permit. The number of food permits and merchandise licenses are limited by law: 5,100 permits and 853 general merchandise licenses are available. Mobile food vending licenses are not limited. Note that a mobile food vending permit is issued to an individual or business to allow for the sale of food from a specific mobile food vending unit, such as a cart or truck. A mobile food vending license authorizes an individual to prepare or serve food from a permitted mobile food vending unit.
There are several versions of a mobile food vendor permit with varying restrictions. Of the 5,100 permits available, 200 are borough permits that limit vendors to one of the five boroughs; 100 are reserved for veterans or vendors with a disability; 1,000 are seasonal and valid only from April to October; 1,000 are green cart permits that limit vendors to selling fruit, vegetables, plain nuts, and water; and 2,800 are unrestricted. Multiple individuals with food vendor licenses can legally operate from one permitted food vending unit. Merchandise licenses are renewed annually, while food licenses and permits are renewed biennially.
There are three relevant exceptions to these rules. The first exception is that merchandise vendors who are veterans are not subject to the limit of 853 licenses. According to data obtained by Mosher and Turnquist (2024), there are approximately 1,000 licensed merchandise vendors who are veterans in New York City. We assume the number of veterans selling merchandise without a license is negligible.
The second exception is First Amendment vendors. A First Amendment vendor sells expressive merchandise such as newspapers, books, and art. Expressive merchandise is considered free speech and protected under the First Amendment of the U.S. Constitution so that the number of First Amendment vendors cannot be restricted by law.
The final exception are concessionaires that operate on New York City parkland through an NYC Parks license or permit. According to the New York City Department of Parks and Recreation, there are approximately 400 concessions within New York City parks. Many offer food services, ranging from food carts to restaurants. We do not consider First Amendment vendors or park concessionaires in this paper. Our estimates are limited to food and merchandise vendors operating outside of parks.
2.2 Survey Design
The Street Vendor Project (SVP), part of the Urban Justice Center, is a non-profit organization that advocates on behalf of New York City street vendors. SVP administered a survey to approximately 2,000 street vendors while distributing Coronavirus relief aid in 2021. The aid was a one-time payment of \$1,000, available to any individual that owned, operated, or was otherwise employed by a street vending business in New York City between 2020 and 2021. First Amendment vendors were also eligible. There were no limits based on the size of the business, the number of sales, or whether licensed and/or permitted. All individuals eligible for aid were invited to complete a survey.
The population of street vendors estimated in this paper is the population eligible for aid (as determined by SVP) that self-identifies as either a food or merchandise vendor. SVP found aid-eligible individuals through membership lists, referrals, and canvassing operations in which SVP affiliates visited neighborhoods. Survey operations continued until \$2,415,000 in funds were exhausted, yielding 2,060 responses.
The survey included 100 questions and was conducted in nine different languages: Arabic, Bangla, Cantonese, English, French, Mandarin, Spanish, Tibetan, and Wolof. The survey items solicited a variety of information from vendors, such as logistical information (e.g., vending location, residential location, and frequency of operation), economic information (e.g., items sold, income, and expenses), and demographic information (e.g., age, race, ethnicity, and immigration status). Most relevant is the fact that respondents classified themselves by the goods they sold (e.g., food vendors, merchandise vendors, First Amendment vendors, etc.), and respondents indicated whether they had the relevant licenses and permits to vend. Of the 2,060 responses, 1,400 identified as food vendors and 559 as merchandise vendors. The remaining 101 respondents were predominantly First Amendment Vendors, which we exclude from our analysis.
Of the 1,400 food vendors, 349 (25%) indicated they had a permit to vend. Of the 559 merchandise vendors, 505 were not veterans, of which 308 (61%) indicated they had a license to vend. The number of respondents is listed by neighborhood in Table 1. (Neighborhoods with few respondents are grouped together.) A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Figure 1. Approximately five percent of vendors did not identify their vending location. These vendors are included in the New York City total, but they are excluded from the neighborhood estimates in Table 1 and Figure 1. (For this reason, the Respondent and Population columns do not sum exactly to the New York City total.) Veteran merchandise vendors are excluded entirely from Table 1 and Figure 1, although these vendors are reflected in our overall estimate of 23,000 vendors.
3. Methodology
We estimate the number of street vendors that operate in New York City. We use ratio estimation, which leverages the fact that the number of licenses and permits are limited by local law. In Section 3.1, we provide a simple explanation of ratio estimation, which is intended to be accessible to a general readership and highlights our main assumption. In Section 3.2, we develop a theoretical justification based on the theory of point processes, which, while more technical, allows us to explicitly state the estimands and underlying assumptions and to derive the standard errors. Inference is conducted using increasing domain asymptotics, and the details are outlined in Sections 7.1, 7.2, and 7.3.
3.1 Ratio Estimation
Ratio estimation is a common approach for estimating the size of a population. See Cochran (1978) and Hald (1998, chap. 16) for a historical discussion and Lohr (2021, chap. 4) for an introduction. We provide a simple explanation based on cross-multiplication, also called the rule of three. The purpose is to highlight the main assumption. A more rigorous derivation is presented in Section 3.2.
Consider a fixed region $A$, and let $N_{i}(A)$ denote the number of individuals in region $A$ of status $i$. To fix ideas, let $A$ be New York City, let $i=0$ indicate mobile food vendors without a permit, and let $i=1$ indicate mobile food vendors with a permit. Also, let $n_{i}(A)$ denote the number of respondents of status $i$.
We observe $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$ from which we would like to determine $N_{0}(A)$, the number of vendors without a permit. The total number of mobile food vendors with or without a permit is then $N_{0}(A)+N_{1}(A)$.
We define the ratio estimator as
$\hat{\Lambda}_{0}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
It is called the ratio estimator because it depends on the ratio of random variables $n_{0}(A)\,/\,n_{1}(A)$. (The total number of permitted vendors, $N_{1}(A)$, is considered fixed.) The estimator is approximately equal to $N_{0}(A)$ when the response rate, $p$, is approximately the same for each status, i.e.,
$p\approx\frac{n_{0}(A)}{N_{0}(A)}\approx\frac{n_{1}(A)}{N_{1}(A)}$
in which case $\hat{\Lambda}_{0}(A)$ is the solution for $N_{0}(A)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$, obtained by cross-multiplication. It follows that $\hat{\tau}(A)=\hat{\Lambda}_{0}(A)+N_{1}(A)$ is approximately equal to the total $N_{0}(A)+N_{1}(A)$.
The main assumption of ratio estimation is that the survey data are representative in the sense that the response rate does not depend on whether a mobile food vendor has a permit. If we further assume that the response rate does not vary by subregion, we can also determine the total number of mobile food vendors by subregion. For example, suppose the response rate in subregion $B\subset A$ is also approximately equal to $p$, i.e.,
$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{0}(B)}{N_{0}(B)}$
Then the subregion estimator
$\tilde{\Lambda}_{0}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
is the solution for $N_{0}(B)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(B)$.
The subregion estimator is useful when $N_{1}(B)$ is not observed and $\hat{\Lambda}_{0}(B)$ cannot be calculated directly. This is the case with the survey data described in Section 2.2. We observe $N_{1}(A)$, the number of permits city-wide, but not $N_{1}(B)$, the number of permits within subregion $B$. (Recall that the vending locations of the respondents are recorded and thus $n_{0}(B)$ and $n_{1}(B)$ are observed for any $B\subset A$.)
The subregion estimator $\tilde{\Lambda}_{0}(B)$ works by replacing the term $N_{1}(B)$ in the ratio estimator $\hat{\Lambda}_{0}(B)$ with
$\dfrac{N_{1}(A)\,n_{1}(B)}{n_{1}(A)}$
which is justified when
$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{1}(B)}{N_{1}(B)}$
This suggests the estimator
$\tilde{\tau}(B)=\tilde{\Lambda}_{0}(B)+\dfrac{N_{1}(A)\,n_{1}(B)}{n_{1}(A)}=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}$
is approximately equal to the subregion total $N_{0}(B)+N_{1}(B)$.
3.2 Model
We provide a point process justification of the ratio estimation procedure described in Section 3.1. The purpose is to explicitly state the estimands and underlying assumptions and to derive the standard errors.
Let $\Pi_{i}$ denote an inhomogeneous Poisson process referencing the location $x\in A\subset\mathbb{R}^{2}$ of each vendor with status $i$ at the time the survey was conducted. As in Section 3.1, we fix ideas by letting $i=0$ indicate mobile food vendors without a permit and $i=1$ indicate mobile food vendors with a permit.
The Poisson process assumption may be justified by the law of rare events. Any sufficiently large area can be partitioned into a large number of theoretically vendable locations. Whether a vendor is located within a partition has a vanishingly small probability, such that the number of vendors in that area is well approximated by a Poisson distribution. This approximation is accurate even under weak dependence (Freedman 1974). Other justifications are possible. For example, see Section 4.2 for a justification of the Poisson process based on the theory of birth processes.
[…]
The data are the locations of the respondents by status (i.e., the realizations of $\pi_{i}$) and the number of vendors in region $A$ with a permit, $N_{1}(A)$. In other words, we assume $n_{0}(B)$ and $n_{1}(B)$ are observed for any $B\subseteq A$, while $N_{0}(B)$ and $N_{1}(B)$ are unobserved with the exception of $N_{1}(A)$. Our goal is to estimate the expected number of street vendors in subregion $B\subseteq A$ without a permit,
[…]
and the expected total with or without a permit,
$\tau(B)=\Lambda_{0}(B)+\mathbb{E}\left[N_{1}(B)\,|\,N_{1}(A)\right]$
We condition on $N_{1}(A)$ to reflect the fact that while the number of vendors of either status in any subregion $B\subsetneq A$ is random, the total number of vendors with permits citywide is fixed by law.
In Section 3.1, we described the main assumption of ratio estimation: that the status and spatial distribution of the respondents is representative. We now make this statement precise. We assume that $p_{i}(x)$ can be decomposed into a constant that does not depend on $i$ or $x$ plus a spatially varying error term, $p_{i}(x)=p\,+\,\epsilon_{i}(x)$, and the error term is orthogonal to the corresponding intensity. i.e.,
[…]
It follows that
[…]
so that by conditioning on $N_{1}(A)$, we arrive at the following probability model for the number of respondents with and without permits
$n_{0}(B)\sim$ $\ \text{Poisson}\big{(}p\,\Lambda_{0}(B)\big{)}$
$n_{1}(B)\,|\,N_{1}(A)\sim$ $\ \text{Binomial}\big{(}N_{1}(B),\,p\,q(B)\big{)}$
where
$q(B)=\frac{\Lambda_{1}(B)}{\Lambda_{1}(A)}$
The maximum likelihood estimates for $\Lambda_{0}(A)$ and $p$ are
$\hat{\Lambda}_{0}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\qquad\text{and}\qquad\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}$
Using increasing domain asymptotics, $\hat{\Lambda}_{0}(A)$ is asymptotically normal with mean $\Lambda_{0}(A)$ and standard error
$\hat{\text{SE}}[\hat{\Lambda}_{0}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
The maximum likelihood estimate for the total $\tau(A)$ is $\hat{\tau}(A)=\hat{\Lambda}_{0}(A)+N_{1}(A)$, which has the same standard error as $\hat{\Lambda}_{0}(A)$. See Section 7.1.2 for details.
The maximum likelihood estimate for $\Lambda_{0}(B)$ is
$\tilde{\Lambda}_{0}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
which is asymptotically normal with mean $\Lambda_{0}(B)$ and standard error
$\tilde{\text{SE}}[\tilde{\Lambda}_{0}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
Finally, the maximum likelihood estimate for the subregion total $\tau(B)$ is
$\tilde{\tau}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}$
The asymptotic standard error is
$\tilde{\text{SE}}[\tilde{\tau}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\frac{1}{n_{1}(A)}-\frac{1}{N_{1}(A)}+\dfrac{n_{1}(B)\,n_{1}(A\backslash B)}{n_{1}(A)\,n_{0}(B)^{2}}}$
where $A\backslash B$ denotes the complement of subregion $B$. See Section 7.1.3 for details.
4. Model Validation
We make two assumptions in Section 3. The main assumption is that the status and spatial distribution of the respondents is representative. This assumption is necessary to ensure the ratio estimator is consistent. A secondary assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous Poisson processes. This assumption is necessary to derive the standard errors.
In this section, we relax both assumptions. Since neither can be relaxed from the data alone, we consider additional information. In Section 4.1, we examine a generalization of the ratio estimator in which sampling weights adjust for differential nonresponse. In Section 4.2, we model the arrival of street vendors as a pure birth process, and we derive a more general formula for the standard errors in which the amount of extra-Poisson variation can be determined by the number of vendors that cluster within markets.
4.1 Representativeness
The main assumption in Section 3 is that the status and spatial distribution of the respondents is representative. If this assumption holds, then the ratio estimator is consistent for a wider class of point processes in which Campbell’s Theorem applies, not just the Poisson process. See Daley and Vere-Jones (2007) for a general statement of Campbell’s Theorem. If this assumption is violated, then $\hat{\Lambda}_{0}(B)$ may be inconsistent. For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(B)\,/\,n_{1}(B)$ would not be representative, and $\hat{\Lambda}_{0}(B)$ would likely underestimate $\Lambda_{0}(B)$, even in large samples.
We think the representativeness assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid as discussed in Section 2. Nevertheless, to evaluate how the estimates might change when this assumption is violated, we generalize the estimator in Section 3.2 to accommodate spatially varying sampling weights. The generalization is useful for determining the sensitivity of our estimates to differential nonresponse as we demonstrate in Section 5.1.
Let $w_{0}(x)$ and $w_{1}(x)$ denote two spatially varying weight functions. We define the weighted counts
$n^{w}_{0}(B)=\sum_{x\,\in\,\pi_{0}\cap B}w_{0}(x)\qquad\text{and}\qquad n^{w}_{1}(B)=\sum_{x\,\in\,\pi_{1}\cap B}w_{1}(x)$
where $\pi_{i}\cap B$ refers to the points of the process $\pi_{i}$ that fall in subregion $B\subseteq A$. We also define the weighted estimators for $\Lambda_{0}(A)$ and $p$,
$\hat{\Lambda}_{0}^{w}(A)=\dfrac{N_{1}(A)\,n^{w}_{0}(A)}{n^{w}_{1}(A)}\qquad\text{and}\qquad\hat{p}^{w}=\frac{n^{w}_{1}(A)}{N_{1}(A)}$
Under Campbell’s Theorem, the means $\mathbb{E}[n^{w}_{0}(A)]$ and $\mathbb{E}[n^{w}_{1}(A)|N_{1}(A)]$ are
[…]
We assume the weight functions $w_{0}$ and $w_{1}$ can be chosen in such a way that the expected value is inversely proportional with the probability of a response. That is,
$p=\mathbb{E}[w_{0}(x)]\,p_{0}(x)=\mathbb{E}[w_{1}(x)]\,p_{1}(x)$
It follows that $\mathbb{E}[n^{w}_{0}(A)]$ and $\mathbb{E}[n^{w}_{1}(A)\,|\,N_{1}(A)]$ simplify to
$\mathbb{E}\left[n^{w}_{0}(A)\right]=p\,\Lambda_{0}(A)\qquad\text{and}\qquad\mathbb{E}\left[n^{w}_{1}(A_{i})\,|\,N_{1}(A)\right]=p\,N_{1}(A)$
and that $\hat{\Lambda}_{0}^{w}(A)$ is approximately normal with mean $\Lambda_{0}(A)$ and standard error
$\hat{\text{SE}}[\hat{\Lambda}_{0}^{w}(A)]=\dfrac{N_{1}(A)\,n_{0}^{w}(A)}{n_{1}^{w}(A)}\sqrt{\dfrac{\hat{\mu}^{2}_{0}(A)+\hat{\sigma}_{00}(A,A)}{n_{0}^{w}(A)^{2}}+\dfrac{\hat{\mu}_{1}^{2}(A)}{n^{w}_{1}(A)^{2}}+\dfrac{(N_{1}(A)-1)\,\hat{\sigma}_{11}(A,A)}{N_{1}(A)\,n^{w}_{1}(A)^{2}}-\dfrac{1}{N_{1}(A)}-\dfrac{2\,\hat{\sigma}_{01}(A,A)}{n_{0}^{w}(A)\,n_{1}^{w}(A)}}$
where
$\hat{\mu}_{k}^{2}(A)=\sum_{x\,\in\,\pi_{k}\cap A}\mathbb{E}\bigl{[}w_{k}^{2}(x)\bigr{]}\qquad\text{and}\qquad\hat{\sigma}_{kl}(A,A)=\sum_{\begin{subarray}{c}x\,\in\,\pi_{k}\cap A\\[1.39998pt] y\,\in\,\pi_{l}\cap A\\[1.39998pt] x\neq y\end{subarray}}\mathbb{C}\bigl{(}w_{k}(x),\,w_{l}(y)\bigr{)}$
See Section 7.1.4 for details including derivations for the subregion and subtotal estimators.
The standard error derived in this section for $\hat{\Lambda}_{0}(A)$ is identical to the standard error $\hat{\Lambda}_{0}(A)$ from Section 3.2 if $w_{0}(x)=w_{1}(x)=1$. In this case, $\hat{\mu}_{0}^{2}(A)=n_{0}(A)=n_{0}^{w}(A)$, $\hat{\mu}_{1}^{2}(A)=n_{1}(A)=n^{w}_{1}(A)$, and $\hat{\sigma}_{00}=\hat{\sigma}_{01}=\hat{\sigma}_{11}=0$.
Note that $\hat{\Lambda}_{0}(A)$ and $\hat{\Lambda}_{0}^{w}(A)$ are related by the relationship
$\hat{\Lambda}_{0}^{w}(A)=\frac{N_{1}(A)\ n^{w}_{0}(A)}{n^{w}_{1}(A)}=\hat{\Lambda}_{0}(A)\,\frac{n^{w}_{0}(A)\,/\,n_{0}(A)}{n^{w}_{1}(A)\,/\,n_{1}(A)}$
It follows that the relative bias depends only on the ratio of the average weights $n^{w}_{0}(A)\,/\,n_{0}(A)$ and $n^{w}_{1}(A)\,/\,n_{1}(A)$. For example, if the average permitted vendor is twice as likely to respond as the average unpermitted vendors, $n^{w}_{0}(A)\,/\,n_{0}(A)$ is twice the size of $n^{w}_{1}(A)\,/\,n_{1}(A)$, and $\hat{\Lambda}_{0}(A)$ is twice the size of $\hat{\Lambda}_{0}^{w}(A)$.
The estimators $\hat{\tau}(A)$ and $\hat{\tau}^{w}(A)$ are similarly related by the relationship
$\hat{\tau}^{w}(A)=\frac{N_{1}(A)\left(n^{w}_{0}(A)+n^{w}_{1}(A)\right)}{n^{w}_{1}(A)}=\hat{\tau}(A)\,\frac{\left(n^{w}_{0}(A)+n^{w}_{1}(A)\right)\,/\,\left(n_{0}(A)+n_{1}(A)\right)}{n^{w}_{1}(A)\,/\,n_{1}(A)}$
The same is true for subregion and subtotal estimators, see Section 7.1.4.
4.2 Overdispersion
The second assumption in Section 3.2 is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes. We consider this assumption secondary because $\hat{\Lambda}_{0}(A)$ is consistent under the wide class of point processes described in Section 4.1. That said, the standard errors derived in Section 3.2 do not necessarily hold when the Poisson assumption is violated.
We think the Poisson assumption is reasonable because a wide class of data generating processes are well-approximated by the Poisson process under the law of rare events as discussed in Section 3.2. Nevertheless, to evaluate how the standard errors might change when this assumption is violated, we consider a general model that describes how street vendors arrived at their present locations. Specifically, we consider two families of pure birth processes: one with a constant arrival rate (or “birth rate”) in which the Poisson assumption holds and one with a linear arrival rate in which the Poisson assumption does not hold.
Suppose vendors serve customers who congregate at fixed locations, which we call markets. Let $M(B)$ denote the number of markets in any subregion $B\subseteq A\subset\mathbb{R}^{2}$. If every market is in a competitive equilibrium such that the expected profit for a new vendor is the same at each market, then the time between vendor arrivals at each market may be well-approximated by an exponential distribution with a constant arrival rate $\theta_{i}$. Further suppose each market starts with one vendor at $t=0$. Then at the time the survey was conducted ($t=1$), the number of vendors, $N_{i}(B)$, would follow a Poisson distribution with mean $\Lambda_{i}(B)=\theta_{i}\,M(B)$, and the results of Section 3.2 hold.
Now suppose instead that larger markets grow faster than smaller markets, an empirical phenomenon known as preferential attachment. Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate, or Yule process. That is, let $N^{t}_{i}(B)$ denote the number of vendors in subregion $B$ at time $t$ so that the time between vendor arrivals at each market follows an exponential distribution with rate $\theta_{i}\,N^{t}_{i}(B)$. If we assume further that each market started with one vendor at $t=0$, then the number of vendors at the time the survey was conducted ($t=1$), $N_{i}(B)$, would follow a negative binomial distribution with mean $\Lambda_{i}(B)=\exp(\theta_{i})\,M(B)$ and variance $\Lambda_{i}(B)\bigr{(}\Lambda_{i}(B)-M(B)\bigr{)}\,/\,M(B)$. i.e.,
$N_{i}(B)\sim\ \text{Negative Binomial}\left(\Lambda_{i}(B),\ \Lambda_{i}(B)\,\frac{\Lambda_{i}(B)-M(B)}{M(B)}\right)$
where we parameterize the negative binomial distribution by its mean and variance for comparison with Section 3.2. In this case, $N_{i}(B)$ exhibits extra-Poisson variation, and the results of Section 3.2 do not hold. See Ross (2014, chap. 6) for details.
To derive the standard errors, suppose a portion of the markets, $p$, were selected for the survey at random so that
$n_{i}(B)\sim\ \text{Negative Binomial}\left(p\,\Lambda_{i}(B),\ p\,\Lambda_{i}(B)\,\frac{\Lambda_{i}(B)-M(B)}{M(B)}\right)$
The conditional distribution of $n_{1}(B)$ given $N_{1}(A)$ is then
$n_{1}(B)\,|\,N_{1}(A)\sim\ \text{Negative Hypergeometric}\bigl{(}N_{1}(A)-1,\ M(A)-1,\ p\,q(B)\,M(A)\bigl{)}$
The conditional mean and variance of $n_{1}(A)$ given $N_{1}(A)$ is
$\mathbb{E}\left[n_{1}(A)\,|\,N_{1}(A)\right]=p\,N_{1}(A)\qquad\text{and}\qquad\mathbb{V}\bigr{(}n_{1}(A)\,|\,N_{1}(A)\bigr{)}=u_{1}(A)\,p\,(1-p)\,N_{1}(A)$
so that under increasing domain asymptotics, the ratio estimator
$\hat{\Lambda}_{0}(A)=\frac{N_{1}(A)\ n_{0}(A)}{n_{1}(A)}$
is asymptotically normal with mean $\Lambda_{0}(A)$ and standard error
$\hat{\text{SE}}[\hat{\Lambda}_{0}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\hat{u}_{0}(A)\left(\dfrac{1}{n_{0}(A)}\right)+u_{1}(A)\left(\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}\right)}$
where
$\hat{u}_{0}(B)=\cfrac{\dfrac{N_{1}(A)\ n_{0}(B)}{n_{1}(A)}-M(B)}{M(B)}\qquad\text{and}\qquad u_{1}(A)=\dfrac{N_{1}(A)-M(A)}{M(A)+1}$
See Section 7.1.5 for details including derivations for the subregion and subtotal estimators. Note that our parameterization makes use of the relationship $q(B)=M(B)\,/\,M(A)$.
The standard error derived in this section for $\hat{\Lambda}_{0}(A)$ is identical to the standard error for $\hat{\Lambda}_{0}(A)$ from Section 3.2 except for the weights $u_{0}$ and $u_{1}$. These weights are the result of the extra-Poisson variation induced by preferential attachment—that is, the linear “birth” rate in which the arrival of new vendors is proportional to the number of vendors in the market. For example, if there are five vendors in each market on average, then $u_{0}$ and $u_{1}$ are approximately 4 and the standard error of the ratio estimator derived in this section is twice the size of the standard error derived in Section 3.2. The same is true for subregion and subtotal estimators, see Section 7.1.5.
5. Data Analysis
We apply the results presented in Section 3 and 4 to the data described in Section 2. In Section 5.1, we use the ratio estimator described in Section 3.1 and 3.2 to estimate the number of street vendors in New York City. In Section 5.2, we use the extensions described in Section 4.1 and 4.2 to deteremine the sensitivity of the estimate to the underlying assumptions. Sensitivity is assessed using additional data from the NYC Office of Administrative Trials and Hearings (OATH) (accessed 2024-04-03) and the American Community Survey from the U.S. Census Bureau (accessed 2022-12-22).
5.1 Estimation
According to the survey data described in Section 2.2, $n_{1}(A)=349$ mobile food vendor respondents indicated they had a permit, and $n_{0}(A)=1,051$ indicated they did not. Since there are $N_{1}(A)=5,100$ permits available, it follows that the number of vendors without permits is 15,3505,100×1,051/ 349formulae-sequence153505100105134915,350$\approx$ 5,100$\times$ 1,051\,/\,34915 , 350 ≈ 5 , 100 × 1 , 051 / 349. Combined with the $5,100$ permitted vendors, the total number of mobile food vendors is approximately $20,500$.
By similar reasoning—i.e., exploiting the fact that the number of general merchandise licenses is limited for non-veterans—the number of non-veteran merchandise vendors is estimated to be approximately $1,400$. When combined with the estimated $1,000$ veteran merchandise vendors (see Section 2.1) and $20,500$ food vendors, the total number of street vendors is approximately $23,000$.
Estimates for select subregions are listed in Table 1. We find that approximately one quarter of street vendors are located in the Manhattan neighborhoods of Chelsea, Clinton, and Lower Manhattan. Another quarter is located in West and North Queens. We also report the margin of error (i.e., two standard errors, half of the width of a 95% confidence interval).
These estimates can be checked by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of street vendors was documented through on-the-ground observations. The independent assessment identified 188 total street vendors near Fordham Road (Fordham Road BID and Street Vendor Project 2024). According to the survey data, 17 respondents indicated that they operate near Fordham Road, yielding an estimate of 166 street vendors total with a standard error of 54. We conclude this estimate is consistent with the 188 street vendors counted independently by Fordham Road BID and Street Vendor Project (2024).
{ThreePartTable}{TableNotes}
This table provides estimates for the number of mobile food vendors and non-veteran general merchandise vendors. The number of veteran general merchandise vendors is approximately 1,000. First amendment vendors and vendors in NYC Parks are not included.
Table 1: Number of Street Vendors
Respondents Population Margin of Error
Bronx
   Bronx Park and Fordham 68 757 239
   Southeast Bronx 28 385 160
   High Bridge and Morrisania 20 233 127
   Central Bronx 9 96 81
   Other Bronx 31 264 137
   Total 156 1,735 376
Brooklyn
      Northwest Brooklyn 46 613 204
      Bushwick and Williamsburg 41 433 175
      Sunset Park 26 344 153
                  Flatbush 19 242 128
            Borough Park 10 134 94
            Southwest Brooklyn 6 64 66
            Other Brooklyn 28 374 159
            Total 176 2,205 418
Manhattan
            Chelsea and Clinton 341 3,289 490
         Lower Manhattan 105 1,250 287
         Gramercy Park and Murray Hill 97 1,062 274
         Lower East Side 65 808 232
         Greenwich Village and Soho 61 726 222
         Upper East Side 56 534 194
         Central Harlem 46 483 185
         Inwood and Washington Heights 33 459 176
         Upper West Side 48 429 172
         East Harlem 38 425 174
         Total 890 9,464 879
Queens
         West Queens 396 4,650 692
      North Queens 67 896 251
      Northwest Queens 23 312 144
      Jamaica 14 181 110
      Other Queens 31 406 168
      Total 531 6,445 860
Staten Island
      Total 4 58 58
New York City
   Total 1,905 21,857 1,941
[…]
Figure 1: Location of Street Vendors
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image]
These maps compare the spatial distribution of respondents (left) to the distribution of vendors suggested by administrative data (New York City Office of Administrative Trials and Hearings (OATH), top right) and federal data (American Community Survey (ACS), bottom right).
5.2 Model Validation
We make two assumptions in Section 3. The main assumption is that the status and spatial distribution of the respondents is representative. We believe trust in SVP and its mission, along with the aid, removed many barriers that typically prevent survey operations from reaching hard-to-reach populations. Instead, nonresponse reflects chance variation in the personal circumstances of a vendor that are largely unrelated to location.
Nevertheless, it is possible that canvassing operations systematically missed locations. We thus compare the spatial distribution of vending locations reported by respondents to the spatial distribution of vendors who allegedly violated street vending laws during the year 2021. We consider the location of an individual’s first violation in 2021, reported in administrative data from the NYC Office of Administrative Trials and Hearings (OATH) (accessed 2024-04-03). We also compare the residencies reported by the respondents to the residencies of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers (Standard Occupational Classification 41-9091) estimated in the 2017-2021 American Community Survey (ACS) Public Use Micro Sample from the U.S. Census Bureau (accessed 2022-12-22). These comparisons are visualized in Figure 1. We find the spatial distribution of vending locations largely agree (top panels) as do the spatial distribution of the residencies (bottom panels). This suggests that, if there were a region of street vendors missed by the SVP, it was also missed by law enforcement and the ACS.
We can also combine the data from OATH and ACS with the results from Section 4.1 to assess the sensitivity of the estimates in Section 5.1 to differential nonresponse, as we now demonstrate. While the difference between these new estimates and the estimates provided in Section 5.1 are not statistically significant at the 95% level, we believe they provide insight into how the estimates in Section 5.1 might theoretically err.
First, we consider the scenario in which the probability a vendor responds is inversely proportional to the average number of vendors cited for violating street vending laws in the surrounding ZIP Code each day, as reported in the OATH data for the year 2021. This might be the case if vendors in high enforcement areas are reluctant to interact with survey takers. The weights are the inverse probabilities—the average number of enforcement actions taken against vendors in the surrounding ZIP Code. We find the weighted counts for food vendors are $n^{w}_{0}(A)=25.7$ and $n^{w}_{1}(A)=5.24$. This suggests there are $\left(n_{0}^{w}(A)+n_{1}^{w}(A)\right)\,/\,\left(n_{0}(A)+n_{1}(A)\right)\,/\left(n_{1}^{w}(A)\,/\,n_{1}(A)\right)=47\%$ more food vendors than estimated in Section 5.1. A similar argument for merchandise vendors suggests there are 9% more merchandise vendors than estimated in Section 5.1. Put together, the weights suggest the number of street vendors in New York City is 40% higher than estimated in Section 5.2.
The second scenario we consider is one in which the probability a vendor responds in proportional to the average number of vendors cited for violating street vending laws in the surrounding ZIP Code. This might be the case if vendors in high enforcement areas are easier to identify and thus easier to survey. Under this scenario, there are 8% fewer food vendors than estimated in Section 5.1 and 9% fewer merchandise vendors, suggesting the number of street vendors in New York City could be 8% lower than estimated in Section 5.2.
The final two scenarios are identical to the first two scenarios, except we use the number of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in the surrounding Public Use Microdata Area as estimated from the 2017-2021 U.S. Census Bureau (accessed 2022-12-22). The result is less extreme than using the average number of vendors cited. If the probability of response is inversely proportional to this estimate, there are 33% more vendors New York City than estimated in Section 5.2. If the probability of response is proportional to this estimate, there are 5% fewer vendors. Combining these four scenarios suggests a conservative range of between 21,000 and 32,200 street vendors operate in New York City.
The second assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous spatial Poisson processes. We argue that this assumption is reasonable because a wide class of data generating processes are well-approximated by the Poisson process under the law of rare events as discussed in Section 3.2.
It is possible, however, that vendors cluster within markets as described in Section 4.2. The weights $u_{0}$ and $u_{1}$ in Section 4.2 correspond to the relative number of vendors without permits and with permits per market respectively. These weights cannot be estimated from the survey data alone. However, we believe that $\Lambda_{0}(A)$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$. It follows that $u_{0}$ and $u_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most 3 times larger if the Poisson assumption is violated and the Yule process holds. We suspect this bound is conservative, however, and the typical market contains only one or two vendors.
6. Discussion
We discuss the implications of our estimates. In particular, we assess the completeness of the New York City Office of Administrative Trials and Hearings (OATH) and American Community Survey (ACS) data. Our estimates suggest that while both data sets are spatially representative, they miss the majority of street vendors. We provide several explanations for this discrepancy.
We estimate 23,000 vendors work in New York City. In comparison, the 2017-2021 ACS estimates 4,634 individuals work the occupation of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in New York City. This suggests that, according to the ACS, there are at most 4,634 street vendors in New York City, and thus our finding indicates that the ACS estimate misses 80% of street vendors or more.
A fairer comparison might account for the fact that the standard error of our estimate is approximately 1,000, and the standard error of the ACS estimate is approximately 488, which we obtained from the replicate weights. Using the lower limit of a 95% confidence interval for our estimate (21,000) and the upper 95% confidence limit for the ACS estimate (5,610), our findings suggest the ACS misses roughly three quarters of street vendors or more.
Note the importance of the assumptions stated in Section 3.2 since deviations might suggest the ACS misses a smaller or larger portion of the street vending population. For example, if the average unpermitted or unlicensed vendor is 10 times more likely to respond than the average permitted or licensed vendor, then the ACS estimate may miss less than half of the street vending population. For another example, if the standard errors from Section 4.2 hold, then the confidence intervals would overlap if there were more than 300 vendors in the average market. We believe it is highly unlikely that our assumptions are violated to this extreme, suggesting the two estimates are inconsistent even under differential nonresponse and overdispersion.
Additional evidence comes from the OATH data, which indicate approximately 2,500 (unique) vendors are summoned to court each year for violating vending laws. If the ACS estimate holds, then more than half of vendors are summoned to court each year. If our estimate holds, then approximately one in ten vendors are summoned to court each year. We believe the latter is more realistic.
There may be several reasons why our estimates are significantly larger than the ACS. The fact that the Street Vendor Project distributed relief aid may have encouraged respondents who are unlikely to respond to the ACS. Moreover, the Street Vendor Project, through membership lists, referrals, and canvassing may have covered individuals whose households were missed by the Census Bureau’s sampling frame. Indeed, a street vendor must be sufficiently visible while on the street in order to sell their items, while the residency of that same vendor may be more easily overlooked.
Another explanation is that the Street Vendor Project survey and the ACS may be measuring different populations or concepts. For example, the ACS estimate reflects the average number of vendors between 2017 and 2021, while our estimate reflects 2020 and 2021. For another example, the Census Bureau determines the occupation of an ACS respondent by autocoding responses to write-in questions. It is possible that the Street Vendor Project’s definition of a vendor is more inclusive. Street vending often provides supplemental income, and thus the discrepancy may reflect the broader challenge of studying the “gig economy.” In this case, either estimate may be relevant, depending on the intended use case.
We believe this application highlights the value of modeling respondents as a point process. We believe our approach is relevant for a wide variety of surveys in which units cannot be chosen at random from a sampling frame. For example, individuals may be sampled on a beach or at various entrances of a park. Future work might explore additional point process models that describe the behavior of the population. For example, we have found that when vendors cluster within markets, standard errors become overdispersed. However, it is possible that vendors protect their turf, leading to underdispersion. It may even be that some vendors benefit from operating near each other, while others benefit from maintaining their distance.